OpenStudy (anonymous):

Exact values of reference angles? sin 9pi/4 tan 19pi/4 sec -3pi/4

OpenStudy (luigi0210):

Reference angle is the angle closest to the x-axis.

OpenStudy (campbell_st):

well you need to decide which quadrant you're in \[2\pi = 1 \times revolution\] so \[\sin(\frac{9\pi}{4}) = \sin(2\pi + \frac{\pi}{4}) = \sin(\frac{\pi}{4})\] which is an exact value \[\tan(\frac{19\pi}{4}) = \tan(4\pi + \frac{3\pi}{4}) = \tan(\frac{3\pi}{4})\] which is an exact value in the 2nd quadrant where tan is negative the last 1 you move backwards by the angle... so its 3rd quadrant... where sec is negative hope it helps

OpenStudy (luigi0210):

At the x-axis, the radians there are: 

OpenStudy (anonymous):

Thanks both of you, that really helped!